OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (agent0smith):

Multiply all the corresponding components together, and add up the results. eg if v = <5, 3, -2> and w = <1, 0, 2> then v * w = 5*1 + 3*0 + (-2)*2 = 1

OpenStudy (anonymous):

@agent0smith so do i do that ^^ but for all 3?

OpenStudy (agent0smith):

You don't need to. You only need v*w, you can ignore r!

OpenStudy (anonymous):

\[v.w=\left( 2 \right)\left( -2 \right)+\left( -8 \right)\left( 6 \right)+\left( -8 \right)\left( -5 \right)\]

OpenStudy (anonymous):

@agent0smith may i ask why is r there? i'm sorry this chapter has confused me

OpenStudy (agent0smith):

v= <2, -8, -8> w= <-2, 6, 5> v ⋅ w can be found by following the example above.

OpenStudy (agent0smith):

r is there because it's there :P

OpenStudy (anonymous):

@agent0smith haha thank you ♥

OpenStudy (agent0smith):

r is just a vector which you do not need. Don't get confused by unnecessary information.

OpenStudy (anonymous):

@agent0smith Okay thank you so much c:

OpenStudy (anonymous):

@agent0smith so for mine it would be : v*w= (2)(-2)+ (-8)(6)+(-8)(-5)= 12

OpenStudy (agent0smith):

Yes! :)